Grade 12Electric Current

a battery of 10 volt and negligible internal resistance is connected to the diagonally opposite corners of a cube network consisting of 12 resistors each of resistance 2 ohm determine the equivalent resistance of the network

Riya rathi
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
 
efective across the diagonally opposite corners of a cubical network consisting of 12 resistors each of resistanvce 2 ohm = 5/6 ohm
Then current through battery = V/R = 10/(5/6) = 12 A